OpenStudy (anonymous):

Solve the system of equations by graphing then classify the system as consistent or inconsistent and as dependent or independent 7x - 9y = 18 9y - 7x = -18

OpenStudy (anonymous):

I can answer you in a minute :-)

OpenStudy (anonymous):

ok, you need to put EACH equation into slope-intercept form.....gosh this is a perfect example of solving by addtion!

OpenStudy (anonymous):

ok, -9y= -7x+18 and 9y = 7x -18 now divide both sides by -9 (first equation) to get y=7/9 x -2 diveide 2nd equ. by a postive 9 tp get y = 7/9 x -2 both are the same. :-) so....

OpenStudy (anonymous):

graph the first equation first by location the y-intercept (-2) then graph out the slope 7/9 so from the -2 (or point (0, -2)) go up 7 and to the right 9....make a dot there and connect it to the (0, -2) This is your graph.....

OpenStudy (anonymous):

YOu there?

OpenStudy (anonymous):

Yes I am following you thank you I am trying to write everything down.

OpenStudy (anonymous):

is the system consistent or inconsistent and dependent or independent?